QUESTION # 2 The following information pertains to Glass Works Company for 2014.

Budgeted direct-labour cost: --------------------------------75,000 hours at $16 per hour

Actual direct-labour cost: ------------------------------------80,000 hours at $17.50 per hour.

Budgeted Manufacturing Overhead: ---------------------$ 997,500

Actual Manufacturing Overhead: Depreciation ------------------------------------------------- $ 240,000

Property Taxes---------------------------------------------------12,000

Indirect Labour --------------------------------------------------82,000

Supervisory Salaries --------------------------------------------200,000

Utilities------------------------------------------------------------ 59,000

Insurance ----------------------------------------------------------30,000

Rental of Space-------------------------------------------------- 300,000

Indirect Material --------------------------------------------------- 79,000

Required a) Compute the firm’s predetermined overhead rate, which is based on direct-labour hours. (3 marks)

b) Calculate the overapplied or underapplied overhead for 2014. (4 marks)

c) Prepare a journal entry to close out the Manufacturing Overhead Account into Cost of Goods Sold. (3 marks)
Answer #1

a) Predetermine overhead rate = 997500/75000 = 13.30 per hour

b) Actual overhead = 1002000

Applied overhead = 80000*13.30 = 1064000

Over applied overhead = 1064000-1002000 = 62000

c) Journal entry

Date account and explanation Debit Credit
manufacturing overhead 62000
Cost of goods sold 62000
(To record over applied overhead)